# AI Agentic Vulnerability Injection And Transformation with Optimized Reasoning

### v3

@@ III-B Design Methodology and Iterative Refinement
Each stages was decomposed into elementary subtasks and mapped to specialized agents. For instance, comprehension was assigned to the Code Understanding and Vuln.-Specific Code Understanding agents, and review to a combination of diff-based agents and static analysis. This decomposition reduces the reasoning burden on individual agents and enables the integration of both LLMs and deterministic tools.
The workflow was then refined through repeated experimentation on the SARD test suite 100 (described in §IV-A), which served as our primary development benchmark. Early prototypes revealed systematic shortcomings: the injector often produced truncated code, over-commented edits, or trivial modifications (e.g., whitespace changes). To address these failure modes, we introduced new agents incrementally. Examples include the Polisher, designed to normalize formatting and repair truncated outputs, and the diff-based refinement loop, which detects trivial edits and triggers corrective re-injection.
Crucially, these refinements targeted generic error patterns rather than dataset-specific artifacts. By focusing on issues such as non-substantive edits and truncation, we ensured that improvements generalized across datasets. As shown later in section V, the same workflow behavior transfers to distinct datasets, confirming that the iterative process captured robust error-recovery patterns rather than overfitting to SARD 100.
@@ III-C Vulnerability Injection Module
The Vulnerability Injection Module converts a benign function $c^{b}$ into a candidate vulnerable variant $c^{v}$ of a specified CWE. It instantiates the first three stages of the analyst-inspired workflow: comprehension, vulnerability-specific analysis, and transformation. The module comprises the following agents (Figure 1):
@@ III-C2 Information Loader (tool)
Examines the benign code to identify elements that prevent the intended vulnerability from manifesting. These include explicit sanitization checks (e.g., bounds validation) as well as broader safety mechanisms such as correct memory management, error handling, or defensive programming patterns. The agent produces a structured inventory of these safeguards, which downstream agents use to determine how the code can be realistically transformed into a vulnerable variant. This process mirrors traditional symbolic static analysis and taint-tracking techniques [staticAnalysis_schwartz2010], but applied inversely. Instead of proving correctness, it highlights the logic that enforces safety. This reverse reasoning is critical to selecting feasible and non-trivial injection points.
III-C4 Vulnerability Injector (LLM + RAG + fine-tuning).
@@
To specialize the vulnerability injection agent without incurring the computational cost of full model fine-tuning, we apply Low-Rank Adaptation (LoRA) [lora_hu2021]. This technique is particularly effective in settings where the base model is large, but the task-specific dataset is limited in size. Rather than updating the full weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, LoRA learns a low-rank decomposition by injecting two trainable matrices $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times k}$, with $r\ll\min(d,k)$, into the model’s architecture: $W=W_{0}+BA$. This reduces the number of trainable parameters while retaining the benefits of large-scale pretraining.
@@
While SFT optimizes for token-level accuracy, it does not directly focus on structural and semantic nuances of code. Therefore, we also apply reinforcement learning (RL) via Group Relative Policy Optimization (GRPO) [grpo_shao2024], using the CodeBLEU [codebleu_ren2020] metric as a reward function.
@@
with $\alpha+\beta+\gamma+\delta=1$. By default, $\alpha=\beta=\gamma=\delta$ [codebleu_ren2020].
@@ III-D1 Difference (tool): Triviality Gate
Invoked only when the triviality gate detects no or trivial edits only. This agent re-attempt to generate $c^{v}$ from scratch, incorporating the context provided by prior agents, with the aim of increasing the likelihood that the agent actually introduces the vulnerability this time. The new candidate is returned to the Difference (Triviality Gate) for re-evaluation. This loop is bounded (default: 10 attempts) to prevent cycling.
@@ III-D8 Vulnerability Fixer (LLM) and Controlled Restart
Attempts a constrained repair using all accumulated context (code, diffs, CWE, rationales, diagnostics). The repaired candidate re-enters validation at the Lightweight Static Analyzer agent up to one time. If the intended vulnerability still does not materialize, the controller restarts from the initial Vulnerability Injector, exploiting LLM stochasticity to explore a different edit trajectory. To ensure termination, the workflow is bounded by a fixed number of correction iterations: 10.
@@ IV Experimental Setup
This section details the experimental setup designed to systematically evaluate the AVIATOR framework. We aim to answer the following research questions (RQs):
@@
  • RQ3: How do different compositions of agents affect AVIATOR’s overall vulnerability generation performance?
IV-A Benchmark Construction
@@
We evaluate AVIATOR specifically for C/C++ code, given the inherent susceptibility of these languages to low-level vulnerabilities (e.g., memory errors), focusing on the function level in alignment with prior vulnerability detection and repair studies [CppUnsafe_erlingsson2025]. We used three complementary popular open-source datasets used in previous studies [sard100_hoole2016] [formai_tihanyi2023] [PrimeVul_ding2024].
@@
  • SARD test suite 100 [nist_sard_2018, sard100_hoole2016]: Contains 102 compact, synthetic test cases, including 34 vulnerable/benign function pairs that can be detected through formal verification tools such as ESBMC [esbmc_gadelha2018]. Their compact and self-contained design, makes them ideal for rapid, fully automated validation and served as our primary development benchmark.
@@
  • FormAI [formai_tihanyi2023]: Includes more realistic and structurally complex synthetic examples. Since it does not provide aligned benign/vulnerable pairs, we manually created benign versions for 37 randomly sampled vulnerable functions, without altering the program’s intent and ensuring no vulnerabilities were detectable by formal verification tools on the resulting benign versions. ESBMC is again used for manual validation, offering a stronger challenge to our injection framework.
@@
  • PrimeVul [PrimeVul_ding2024]: Contains over 5,000 real-world C/C++ code pairs annotated with vulnerability labels. Due to the lack of full context (e.g., missing class definitions, global variables, and interprocedural dependencies), as well as bigger code samples, formal verification tools such as ESBMC are not directly applicable. We use the provided: (1) training split of 3,789 pairs across 112 CWEs to fine-tune the Vulnerability Injector, (2) validation split of 480 pairs (63 CWEs) as the retrieval knowledge base in AVIATOR’s RAG setup, and (3) 45 randomly sampled pairs from the test split (24 CWEs) for manual evaluation of injection success.
@@ IV-B Implementation Details
We experimented with two base large language models for the AI agents in our workflow. Qwen2.5-Coder-32B-Instruct [qwen25coder_hui2024], is a specialized open-source coding LLM with 32 billion parameters. We additionally experimented with Llama-4-Maverick [llama4_meta2025], a general-purpose, mixture-of-experts-based LLM with approximately 400 billion parameters (17 billion active parameters), to explore the impact of model size and specialization on the agents’ reasoning capabilities.
@@
We use gte-Qwen2-1.5B-Instruct as the embedding model for vulnerable/benign code function pairs drawn from the validation split of PrimeVul (§IV-A) that constitutes the RAG knowledge base. At the time of our experiments, this model achieved state-of-the-art scores on the MTEB benchmark for embeddings up to 8K tokens [mteb_huggingface2023]. Since 99% of functions in the PrimeVul training split fall below this length, no chunking was required. We set the retrieval parameter to $k=4$, following the recommended default.
@@ IV-D Evaluation and Metrics
For compatible test cases [nist_sard_2018] and [formai_tihanyi2023], we automatically measured vulnerability injection success using the formal verification tool ESBMC v7.5 (Efficient SMT-based Bounded Model Checker), recognized for its effectiveness in detecting vulnerabilities in bounded code contexts [esbmc_gadelha2018, Sw_verification_competion_beyer2023]. It implements state-of-the-art incremental Bounded Model Checking and $k$-induction proof-rule algorithms. Namely, if the solver reaches termination within a bound less than or equal to $k$, we can definitively prove the absence of software errors. We set the bound to 100 and timeout to 5 minutes.
@@
  • Pass@k [passk_chen2021]: Estimates the probability that at least one of $k$ independently sampled outputs is correct. If $c$ is the number of correct outputs among $k$ samples:
@@
For PrimeVul, where formal verification tools such as ESBMC are not directly applicable due to the lack of full context (e.g., missing class definitions and global variables), as well as bigger code samples, we manually validated a sample of 45 injected functions drawn uniformly at random, following prior work [d2adatasetbuilt_zheng2021, DiverseVul_yizheng2023, vinj_nong2024, lessIsEnough_wen2023]. The evaluation was performed by three authors experienced in C/C++, including two senior security researchers. Each injected sample was categorized as (1) weakness present, (2) no weakness, or (3) undecidable—typically due to reliance on external functions or context beyond the injected function. Since we lack complete test cases for most programs, we assess the presence of weaknesses rather than confirmed vulnerabilities (which require demonstrable exploitability). The reported injection success rate reflects the percentage of samples found to contain a weakness, based only on those that could be conclusively assessed.
@@ V Results and Analysis
To evaluate AVIATOR’s effectiveness, we assessed its vulnerability injection success using samples from three benchmarks: SARD100, FormAI, and PrimeVul. Table I shows the average success rates for the full AVIATOR workflow ($W_{13}$) with the SFT injection agent. AVIATOR achieved a 95% success rate on SARD100 and 91% on the more complex FormAI benchmark, indicating strong generalization beyond simple patterns.
For PrimeVul, manual analysis was conducted on 45 injected samples due to the lack of external context needed for automated verification. Out of 34 analyzable cases, 32 contained a verifiable weakness within the function body (94%). The remaining samples involved undefined external symbols, so the analysis focused on confirming the presence of weakness insertion rather than evaluating full vulnerability exploitability.
TABLE I: AVIATOR injection success across datasets: Average Injection Success Rate on 5 runs for SARD100 and FormAI, and on analyzable cases for PrimeVul due to limited context.
@@
Table II compares AVIATOR with other techniques, showing that its success rate on FormAI (91%) and PrimeVul (94%) surpasses existing injection methods, like VinJ and VulGen, which report 69% success rates. AVIATOR’s guided injection process allows for precise, context-aware edits that mimic developer like mistakes, making it effective even in partially defined or real-world settings.
@@
TABLE II: Label accuracy and injection success rates (marked with *) for major vulnerability datasets that adhere to BCD properties (see Section II-A). AVIATOR results are based on automated assessment using the FormAI dataset and manual validation using PrimeVul.
Model / Dataset (*) Accuracy Source
Graph2Edit * [Graph2Edit_yao2021] 13% [vulgen_nong2023]
BigVul [BigVul_jiahao2020] 25% [PrimeVul_ding2024]
AVIATOR $W_{1}$ * 31% FormAI validation
CrossVul [CrossVul_georgios2021] 48% [PrimeVul_ding2024]
Getafix * [Graph2Edit_yao2021] 50% [vulgen_nong2023]
CVEFixes [CVEfixes_guru2021] 52% [PrimeVul_ding2024]
DiverseVul [DiverseVul_yizheng2023] 60% [PrimeVul_ding2024]
VulGen * [vulgen_nong2023] 69% [vulgen_nong2023]
VinJ * [vinj_nong2024] 69% [vinj_nong2024]
AVIATOR $W_{13}$ & SFT * 91% FormAI validation
AVIATOR $W_{13}$ & SFT * 94% PrimeVul validation
@@
Table III reports the average injection success rate ($AISR_{5}$) and its standard deviation across 5 vulnerability injection runs on the same data. On the more complex FormAI dataset, SFT yields significant gains over the base model (91% vs. 85%), while also reducing variance. In contrast, GRPO not only fails to outperform the base model on FormAI (dropping to 84%), but also provides no gain on SARD100. These results indicate that in our setup, GRPO is less effective than SFT, despite using the same training data and more training computation.
@@
TABLE III: AVIATOR $W_{13}$ injection success rate ($AISR_{5}$) across 5 vulnerability injection generation with and without fine-tuning, on FormAI and SARD100.
@@
We also reported Pass@k performance to assess the likelihood of producing a successful injection within $k$ generation attempts. Table IV shows that SFT improves results consistently across all $k$ on both benchmarks. On FormAI, Pass@1 improves by over 5 percentage points (from 84.3% to 89.9%), and SARD100 reaches near-perfect performance even for small $k$. These gains further validate the robustness added by SFT.
It is important to note that, even without fine-tuning, AVIATOR achieves strong baseline performance. This indicates that the agentic workflow itself, through decomposition, correction loops, and validation, is sufficient to generate realistic and diverse vulnerabilities. Fine-tuning provide further gains, but is not strictly necessary for the framework to operate effectively. While baseline performance depends on how well the underlying LLM has been trained on the target language, this property remains especially valuable when extending to programming languages or vulnerability classes where aligned training data is scarce, since AVIATOR can still produce rich vulnerability patterns.
TABLE IV: AVIATOR $W_{13}$ pass@k for 10 vulnerability injection generations with and without fine-tuning, on FormAI and SARD100.
@@
Refer to caption Refer to caption
(a) SARD 100 (b) FormAI
@@
V-D Compute and Cost Estimates

### v4

@@ III-B Design Methodology and Iterative Refinement
Each of these stages was decomposed into elementary subtasks and mapped to specialized agents. For instance, comprehension was assigned to the Code Understanding and Vuln.-Specific Code Understanding agents, and review to a combination of diff-based agents and static analysis. This decomposition reduces the reasoning burden on individual agents and enables the integration of both LLMs and deterministic tools.
The workflow was then refined through repeated experimentation on the SARD test suite 100 (described in §IV-A), which served as our primary development benchmark. Early prototypes highlighted systematic shortcomings: the injector frequently returned truncated code, over-commented edits, or trivial modifications (e.g., whitespace changes). To address these failure modes, we introduced new agents incrementally. Examples include the Polisher, designed to normalize formatting and repair truncated outputs, and the diff-based refinement loop, which detects trivial edits and triggers corrective re-injection.
Crucially, these refinements targeted generic error patterns rather than dataset-specific artifacts. By focusing on issues such as non-substantive edits and truncation, we ensured that improvements generalized across datasets. As shown later in section V, the same workflow behavior transfers to very distinct datasets, confirming that the iterative process captured robust error-recovery patterns rather than overfitting to SARD 100.
@@ III-C Vulnerability Injection Module
The Vulnerability Injection Module converts a benign function $c^{b}$ into a candidate vulnerable variant $c^{v}$ of a specified CWE. It instantiates the first three stages of the analyst-inspired workflow: comprehension, vulnerability-specific analysis, and transformation. The module comprises the following agents (Figure 1). We provide the exact prompt templates used by the major agents in AVIATOR in Appendix -A.
@@ III-C2 Information Loader (tool)
Examines the benign code to identify elements that prevent the intended vulnerability from manifesting. These include explicit sanitization checks (e.g., bounds validation) as well as broader safety mechanisms such as correct memory management, error handling, or defensive programming patterns. The agent produces a structured inventory of these safeguards, which downstream agents use to determine how the code can be realistically transformed into a vulnerable variant. This process mirrors traditional symbolic static analysis and taint-tracking techniques [45], but applied inversely. Instead of proving correctness, it highlights the logic that enforces safety. This reverse reasoning is critical to selecting feasible and non-trivial injection points.
III-C4 Vulnerability Injector (LLM + RAG + fine-tuning)
@@
To specialize the vulnerability injection agent without incurring the computational cost of full model fine-tuning, we apply Low-Rank Adaptation (LoRA) [23]. This technique is particularly effective in settings where the base model is large, but the task-specific dataset is limited in size. Rather than updating the full weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, LoRA learns a low-rank decomposition by injecting two trainable matrices $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times k}$, with $r\ll\min(d,k)$, into the model’s architecture: $W=W_{0}+BA$. This reduces the number of trainable parameters while retaining the benefits of large-scale pretraining.
@@
While SFT optimizes for token-level accuracy, it does not directly focus on structural and semantic nuances of code. Therefore, we also apply reinforcement learning (RL) via Group Relative Policy Optimization (GRPO) [47], using the CodeBLEU [43] metric as a reward function.
@@
with $\alpha+\beta+\gamma+\delta=1$. By default, $\alpha=\beta=\gamma=\delta$ [43].
@@ III-D1 Difference (tool): Triviality Gate
Invoked only when the triviality gate detects no or trivial edits only. This agent re-attempt to generate $c^{v}$ from scratch, incorporating the context provided by prior agents, with the aim of increasing the likelihood that the agent actually introduces the vulnerability this time. The new candidate is returned to the Difference (Triviality Gate) for re-evaluation. This loop is bounded (default: 10 attempts) to prevent inifinite cycling.
@@ III-D8 Vulnerability Fixer (LLM) and Controlled Restart
Attempts a constrained repair using all accumulated context (code, diffs, CWE, rationales, diagnostics). The repaired candidate re-enters validation at the Lightweight Static Analyzer agent up to one time. If the intended vulnerability still does not materialize, the controller restarts from the initial Vulnerability Injector, exploiting LLM stochasticity to explore a different edit trajectory.To ensure termination, we cap the total number of correction attempts (default: 10) across reinjection and refinement loops.
@@ IV Experimental Setup
In this section, we detail the experimental design used to evaluate AVIATOR. We present our research questions, the datasets utilized for both injection validation and downstream tasks, and the implementation details of our framework. We aim to answer the following research questions:
@@
  • RQ3: How do different compositions of agents (from monolithic to full agentic workflow) affect AVIATOR’s vulnerability generation capabilities?
  • RQ4: How does AVIATOR’s injection success rate varies across different vulnerability categories (CWEs)?
  • RQ5: Can AVIATOR-augmented datasets improve the performance of state-of-the-art deep learning-based vulnerability detection models compared to existing augmentation techniques?
IV-A Datasets
@@
We evaluate AVIATOR specifically for C/C++ functions, given the inherent susceptibility of these languages to low-level vulnerabilities (e.g., memory errors), focusing on the function level in alignment with prior vulnerability detection and repair studies [15]. For RQ1 to RQ4 we used three complementary popular open-source datasets used in previous studies [22] [48] [12]. For RQ5, we additionally used [56] [17] [12] following the methodology established in VulScribeR [10]. Table I summarizes the statistics of each dataset in our experiments.
TABLE I: Statistics of the Studied Datasets
Dataset Vulnerable Samples Clean Samples Total Ratio CWE Num Entropy
SARD100 [22] 102 102 204 1:1 23 3.98
FormAI [48] 192,712 117,774 310,486 1:0.6 19 3.61
PrimeVul Train [12] 4,802 169,345 174,147 1:35.3 122 4.76
PrimeVul Validation 593 23,347 23,940 1:39.4 73 5.05
PrimeVul Test 549 24,232 24,781 1:44.1 71 5.09
Devign [56] 10,768 12,024 22,792 1:1.1 NA 4.84
BigVul Train [17] 8,783 142,125 150,908 1:16.2 91 4.94
BigVul Validation 1,038 17,826 18,864 1:17.2 84 4.95
BigVul Test 1,079 17,785 18,864 1:16.5 85 4.72
Entropy
We report an embedding-based entropy score as a proxy for code diversity and richness, complementary to CWE count (which captures label-space coverage). Following [10], we embed each function with CodeBERT, reduce embeddings to three dimensions using principal component analysis, discretize the space into a 10-bin histogram, and compute Shannon entropy over the resulting histogram. Higher entropy indicates a broader variety of code structures and idioms in the dataset, whereas lower entropy reflects more homogeneous or template-like code.
@@
  • SARD test suite 100 [34, 22]: Contains 102 compact, synthetic test cases, including 34 vulnerable/benign function pairs that can be detected through formal verification tools such as ESBMC [20]. Their compact and self-contained design, makes them ideal for rapid, fully automated validation and served as our primary development benchmark.
@@
  • FormAI [48]: Includes more realistic and structurally complex synthetic examples. Since it does not provide aligned benign/vulnerable pairs, we manually created benign versions for 37 randomly sampled vulnerable functions, without altering the program’s intent and ensuring no vulnerabilities were detectable by formal verification tools on the resulting benign versions. ESBMC is again used for automated validation, offering a stronger challenge to our injection framework.
@@
  • PrimeVul [12]: A large dataset of more complex real-world code samples. Due to the lack of full context (e.g., missing class definitions, global variables) and larger code size, formal verification tools like ESBMC are not directly applicable. We use: (1) the paired training split with 3,789 corresponding pairs across 112 CWEs to fine-tune the Vulnerability Injector, (2) a subset of the full training set to train downstream DLVD, (3) the validation split of 480 corresponding pairs (63 CWEs) as the retrieval knowledge base in AVIATOR’s RAG setup, (4) 45 randomly sampled pairs from the test split (24 CWEs) for manual evaluation of injection success, and (5) the full test split (71 CWEs) to evaluate downstream DLVD performance.
  • Devign [56]: Following previous studies [36, 37, 10], we use Devign as one of the two primary training datasets for the downstream tasks in RQ5. It offers a relatively balanced distribution between vulnerable and clean samples, making it suitable for training robust detection models without excessive class imbalance issues.
  • BigVul [17]: A large-scale real-world dataset. Consistent with the methodology in VulScribeR [10], we utilize the testing set of BigVul as an independent test set to evaluate the generalization capability of the vulnerability detection models trained on AVIATOR-augmented data (RQ5).
@@ IV-B Implementation Details
We experimented with two base large language models for the AI agents in our workflow. Qwen2.5-Coder-32B-Instruct [24], is a specialized open-source coding LLM with 32 billion parameters. We additionally experimented with Llama-4-Maverick [32], a general-purpose, mixture-of-experts-based LLM with approximately 400 billion parameters (17 billion active parameters), to explore the impact of model size and specialization on the agents’ reasoning capabilities.
@@
We use gte-Qwen2-1.5B-Instruct as the embedding model for vulnerable/benign code function pairs drawn from the validation split of PrimeVul (§IV-A) that constitutes the RAG knowledge base. At the time of our experiments, this model achieved state-of-the-art scores on the MTEB benchmark for embeddings up to 8K tokens [16]. Since 99% of functions in the PrimeVul training split fall below this length, no chunking was required. We set the retrieval parameter to $k=4$, following the recommended default.
@@ IV-D Evaluation and Metrics
For compatible test cases [34] and [48], we automatically measured vulnerability injection success using the formal verification tool ESBMC v7.5 (Efficient SMT-based Bounded Model Checker), recognized for its effectiveness in detecting vulnerabilities in bounded code contexts [20, 2]. It implements state-of-the-art incremental Bounded Model Checking and $k$-induction proof-rule algorithms. Namely, if the solver reaches termination within a bound less than or equal to $k$, we can definitively prove the absence of software errors. We set the bound to 100 and timeout to 5 minutes.
@@
  • Pass@k [6]: Estimates the probability that at least one of $k$ independently sampled outputs is correct. If $c$ is the number of correct outputs among $k$ samples:
@@
For PrimeVul, where formal verification tools such as ESBMC are not directly applicable due to the lack of full context (e.g., missing class definitions and global variables), as well as bigger code samples, we manually validated a sample of 45 injected functions drawn uniformly at random, following prior work [54, 7, 38, 50]. It is important to note that our total evaluation sample size is consistent with prior similar studies [10, 38, 54]. The evaluation was performed by three authors experienced in C/C++, including two senior security researchers. Each injected sample was categorized as (1) weakness present, (2) no weakness, or (3) unknown, typically due to reliance on external functions or context beyond the injected function. Since we lack complete test cases for most programs, we assess the presence of weaknesses rather than confirmed vulnerabilities (which require demonstrable exploitability). The reported injection success rate reflects the percentage of samples found to contain a weakness, based only on those that could be conclusively assessed.
IV-E Downstream Deep Learning-Based Vulnerability Detection (DLVD) Setup
To answer RQ5, we evaluate the practical utility of AVIATOR by using this workflow to augment training data for downstream vulnerability detection tasks. We adopt the experimental protocol established in recent state-of-the-art studies [10], ensuring a fair comparison against existing augmentation techniques.
IV-E1 Datasets and Augmentation Strategy
We utilized two widely-used benchmarks as seed datasets: Devign [56] and PrimeVul [12], described in Section IV-A. For each dataset, we generated 5,000 new vulnerable samples using AVIATOR and applied specific augmentation strategies to handle class balance, consistently with the methodology in [10]:
  • Devign Augmentation: We added the 5,000 AVIATOR-generated vulnerable samples to the training set. To prevent label shift and maintain the original class distribution, we simultaneously added a proportional number of benign samples. We sourced these benign samples from [36], which have been previously deduplicated across datasets.
  • PrimeVul Augmentation: Due to the extreme class imbalance in PrimeVul, we first down-sampled the benign samples in the training set to reduce the ratio from 1:35 to 1:15, making model training possible. After injecting 5,000 AVIATOR-generated vulnerable instances, we applied Random Oversampling (ROS) exclusively to the original vulnerable samples to achieve a balanced 1:1 class ratio for training, to conduct a fair comparison between the methods [10].
For both datasets, we preserved the original distributions of the test sets to ensure that the evaluation assesses performance against real-world distributions.
IV-E2 Models
We evaluated performance across four distinct deep learning architectures, representing two major families of vulnerability detection models:
  • Graph-based Models: We employed Devign [56] and Reveal [5], which utilize Graph Neural Networks (GNNs) on Code Property Graphs (CPGs) to capture structural semantics.
  • Transformer-based Models: We employed LineVul [19], a BERT-based model optimized for line-level detection, and CodeT5 [49], an encoder-decoder model pre-trained on code, used as a strong baseline in recent benchmarks [12].
Training Configuration: To ensure a rigorous comparison, we adhered to the training settings reported in [10]. For the graph-based models (Devign, Reveal), we trained each model 5 times with different random seeds and reported the results from the run achieving the highest F1-score. For Transformer-based models (LineVul, CodeT5), we trained for 10 epochs and selected the checkpoint that achieved the highest F1-score.
IV-E3 Baselines
We benchmarked AVIATOR against four distinct baselines:
  1. 1.
    No Augmentation (No Aug): Training on the original dataset, to establish a reference.
  2. 2.
    Random Oversampling (ROS): A standard class-balancing technique that duplicates vulnerable samples to mitigate imbalance, shown to be a competitive baseline in prior studies [52].
  3. 3.
    VulGen [37] and VGX [36]: State-of-the-art injection techniques that rely on pattern mining and deep learning-based localization.
  4. 4.
    VulScribeR [10]: The current state-of-the-art downstream-optimized generative augmentation method, leveraging LLMs and retrieval-augmented generation (RAG). VulScribeR explicitly targets downstream DLVD performance rather than injection quality, in contrast to AVIATOR.
@@ V Results and Analysis
To evaluate AVIATOR’s effectiveness, we assessed its vulnerability injection success using samples from three benchmarks: SARD100, FormAI, and PrimeVul. Table II shows the average success rates for the full AVIATOR workflow ($W_{13}$) with the SFT injection agent.
AVIATOR achieved a 95% success rate on SARD100 and 91% on the more complex FormAI benchmark, indicating strong generalization beyond simple patterns. For PrimeVul, manual analysis was conducted on 45 injected samples due to the lack of external context needed for automated verification. Out of 34 analyzable cases, 32 contained a verifiable weakness within the function body (94%). The remaining samples involved undefined external symbols, so the analysis focused on confirming the presence of weakness insertion rather than evaluating full vulnerability exploitability.
TABLE II: AVIATOR injection success across datasets: Average Injection Success Rate on 5 runs for SARD100 and FormAI, and on analyzable cases for PrimeVul due to limited context.
@@
Table III compares AVIATOR with other techniques, showing that its success rate on FormAI (91%) and PrimeVul (94%) surpasses existing injection methods, like VinJ and VulGen, which report 69% success rates. AVIATOR’s guided injection process allows for precise, context-aware edits that mimic developer-like mistakes, making it effective even in partially defined or real-world settings.
TABLE III: Label accuracy and injection success rates (marked with *) on major vulnerability datasets satisfying the BCD properties (Section II-A). Results for AVIATOR are obtained via automated evaluation on SARD100 and FormAI, and manual validation on PrimeVul. The reported score for AVIATOR corresponds to the mean injection success rate across all three evaluation datasets.
Model / Dataset (*) Accuracy Source
Graph2Edit * [53] 13% [37]
BigVul [17] 25% [12]
AVIATOR $W_{1}$ * 31% FormAI validation
CrossVul [35] 48% [12]
Getafix * [53] 50% [37]
CVEFixes [3] 52% [12]
DiverseVul [7] 60% [12]
VulGen * [37] 69% [37]
VinJ * [38] 69% [38]
VulScribeR * [10] 82% [10]
VGX * [36] 90% [36]
AVIATOR $W_{13}$ & SFT * 93% Cross-Dataset Eval.
Table III compares AVIATOR with other techniques, showing that its success rate across all three evaluation benchmarks (93%) surpasses prior methods. Earlier injection methods such as VulGen and VinJ that combine pattern mining with learned localization, report an injection success of 69%, implying that a non-negligible fraction of generated samples do not instantiate the intended weakness. VGX follows a similar pattern-mining paradigm and substantially improves injection success (90%), yet does so under tighter constraints, limited CWE coverage (23 types) and a primary focus on single-statement edits, which restricts the space of vulnerabilities it can realistically generate. In parallel, VulScribeR leverages LLMs and RAG to optimize downstream DLVD performance. However, it does not explicitly prioritize injection correctness and reports a lower estimated injection success (82%). In contrast, AVIATOR’s guided, multi-step injection workflow enables precise, context-aware edits that more closely resemble developer-introduced vulnerabilities, allowing it to remain effective even in partially specified or real-world code settings.
@@
Table IV reports the average injection success rate ($AISR_{5}$) and its standard deviation across 5 vulnerability injection runs on the same data. On the more complex FormAI dataset, SFT yields significant gains over the base model (91% vs. 85%), while also reducing variance. In contrast, GRPO not only fails to outperform the base model on FormAI (dropping to 84%), but also provides no gain on SARD100. A plausible explanation is that, in this setting, GRPO may be less well-matched to the available training signal and reward formulation, making it harder to consistently reinforce high-quality injection behaviors. Overall, these results indicate that SFT provides a more stable and effective optimization strategy for vulnerability injection in AVIATOR.
@@
TABLE IV: AVIATOR $W_{13}$ injection success rate ($AISR_{5}$) across 5 vulnerability injection generation with and without fine-tuning, on FormAI and SARD100.
@@
We also reported Pass@k performance to assess the likelihood of producing a successful injection within $k$ generation attempts. Table V shows that SFT improves results consistently across all $k$ on both benchmarks. On FormAI, Pass@1 improves by over 5 percentage points (from 84.3% to 89.9%), and SARD100 reaches near-perfect performance even for small $k$. These gains further validate the robustness added by SFT.
TABLE V: AVIATOR $W_{13}$ pass@k for 10 vulnerability injection generations with and without fine-tuning, on FormAI and SARD100.
@@
Refer to caption Refer to caption
(a) SARD 100 (b) FormAI
It is important to note that, even without fine-tuning, AVIATOR achieves strong baseline performance. This indicates that the agentic workflow itself, through decomposition, correction loops, and validation, is sufficient to generate realistic and diverse vulnerabilities. Fine-tuning provides further gains, but is not strictly necessary for the framework to operate effectively. While baseline performance depends on how well the underlying LLM has been trained on the target language, this property remains especially valuable when extending to programming languages or vulnerability classes where aligned training data is scarce, since AVIATOR can still produce rich vulnerability patterns.
Learning dynamics
Figure 3, left, shows the SFT training loss, which demonstrates stable convergence within five epochs. In contrast, GRPO (Figure 3, right) achieved rapid early gains in reward, reaching a high CodeBLEU score by step 400. While the reward plateaus thereafter, it remains stable, indicating that GRPO quickly finds a strong policy and maintains it. However, this comes at significantly higher computational cost, requiring 10 days of training compared to only 10 hours for SFT.
Refer to caption Refer to caption
[…]
@@
V-D RQ4. Performance Variability Across Vulnerability Categories
To further characterize the capabilities of AVIATOR, we analyze injection success rates across the CWE categories present in our evaluation benchmarks (SARD100, FormAI, and the PrimeVul subset). As shown in Figure 4, this distribution reflects only the CWEs present in the datasets; the current AVIATOR instantiation supports 133 CWE types in total as tested in V-F.
Figure 4: Injection success rate by CWE ID within the evaluation datasets. Bubble size indicates the number of unique code samples available for each CWE. The annotation denotes samples from SARD100 and FormAI evaluated across 5 repeated runs; their reported success rate is the average over the 5 runs. The color gradient reflects the total volume of injection attempts, accounting for all repetitions. CWEs included in the MITRE Top 25 Most Dangerous Software Weaknesses are highlighted with a bold border.
Across the evaluated benchmarks, AVIATOR achieves a high overall injection success rate of 93%. In particular, the framework performs strongly on vulnerability types listed in the MITRE Top 25 Most Dangerous Software Weaknesses. This suggests that AVIATOR is well-aligned with real-world security priorities and effectively injects vulnerabilities that matter most in practice.
At a finer granularity, AVIATOR performs well on both simple and complex vulnerability injections. It reliably injects vulnerabilities realizable through localized edits (e.g., missing checks or unsafe memory accesses), while also succeeding on categories that require non-trivial reasoning over program state and execution order. For instance, AVIATOR achieves high success rates on CWE-416 (Use After Free, $N=11$), CWE-476 (NULL Pointer Dereference, $N=22$), and CWE-120 (Buffer Copy without Checking Size of Input, $N=12$), which require consistent reasoning about data flow, pointer validity, and memory access constraints. Beyond memory safety, AVIATOR also performs reliably on arithmetic and control-dependent vulnerabilities such as CWE-369 (Divide by Zero, $N=5$), demonstrating versatility across vulnerability classes.
However, figure 4 also reveals performance variability primarily on vulnerability categories that either lack precise injection semantics or require constraint satisfaction beyond the function scope. For instance, AVIATOR fails to inject CWE-391 (Unchecked Error Condition), which yields a 0% success rate. This CWE lacks a precise modern definition and is slated for deprecation, making it difficult to formulate a concrete injection objective. Similarly, CWE-190 (Integer Overflow) poses a greater challenge, achieving a 67% success rate, because successful injection often depends on satisfying strict arithmetic constraints that involve global assumptions or external inputs. These cases reflect inherent limitations of function-scoped injection rather than shortcomings of AVIATOR’s workflow.
Answer to RQ4: AVIATOR achieves a high overall injection success rate of 93% and performs strongly on practically relevant vulnerabilities, including multiple CWEs from the MITRE Top 25. The framework successfully injects non-trivial vulnerabilities that require reasoning about data flow, control flow, and object lifetimes (e.g., CWE-416, CWE-476), while lower success rates remain confined to ambiguous CWEs or those requiring precise arithmetic reasoning beyond function-level context.
V-E Compute and Cost Estimates